OpenStudy (anonymous):

Determine the 2 numbers whose difference is 12 and whose product is a maximum.I came up with (x + 6)^2 - 36 But, what do I do now? It is functions, grade 11. Do I need to explain how I got x + 6)^2 - 36?

OpenStudy (anonymous):

put one number as x then since the difference is 12 the other is x - 12

OpenStudy (anonymous):

or you could use x and x + 12, which is what it looks like you used

OpenStudy (anonymous):

although on the face of it i don't see how you can solve this problem. maybe it means the sum is 12, not the difference

OpenStudy (anonymous):

to me "difference" means what you get when you subtract so the difference between 100 and 112 is 12

OpenStudy (anonymous):

That is exactly what I used. So, then it is Product =(x)(x+12) = x^2 + 12x = x^2 +12x + 36 - 36 = (x + 6)^2 - 36 I did this because that's what we are doing now I'm not sure how from that I'm finding these 2 numbers